if Rs.450 amounts to Rs.540 in 4 years at simple interest, then what sum of money will amount to Rs.637.50 in 5 years at the same rate of interest?
Answer by stanbon(75887)

if Rs.450 amounts to Rs.540 in 4 years at simple interest, then what sum of money will amount to Rs.637.50 in 5 years at the same rate of interest?
-------
I = P*r*t
540 = 450*r*4
-----
r = 0.3
------
637.50 = P*0.3*5
P = Rs425
------------
Cheers,
Stan H.
---------